# When Distance Distracts: Representation Distance Bias in BT-Loss for Reward Models

### v1

@@ Abstract
Reward models are central to Large Language Model (LLM) alignment within the framework of RLHF. The standard objective used in reward modeling is the Bradley-Terry (BT) loss, which learns from pairwise data consisting of a pair of chosen and rejected responses. In this work, we analyze the per-sample gradient of BT-loss and show that its norm scales with two distinct components: (1). the difference in predicted rewards between chosen and rejected responses, which reflects the prediction error, and critically, (2). representation distance between the pair measured in the output space of the final layer. While the first term captures the intended training signal, we show that the second term can significantly impact the update magnitude and misalign learning. Specifically, pairs with small representation distance often receive vanishingly weak updates, even when misranked, while pairs with large distance receive disproportionately strong updates. This leads to gradients from large-distance pairs to overshadow those from small-distance pairs, where fine-grained distinctions are especially important. To overcome this limitation, we propose NormBT, an adaptive pair-wise normalization scheme that balances representation-driven effects and focuses learning signals on prediction error. NormBT is a lightweight, drop-in integration to BT loss with negligible overhead. Across various LLM backbones and datasets, NormBT improves reward model performance consistently, with notable gains of over $5\%$ on the Reasoning category of RewardBench, which contains numerous small-distance pairs. This work reveals a key limitation in the widely used BT objective and provides a simple, effective correction.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as the leading framework for aligning Large Language Models (LLMs) with human preferences (christiano2017deep; ziegler2020finetuninglanguagemodelshuman; ouyang2022traininglanguagemodelsfollow). Instead of traditional supervision, RLHF uses preference signals to guide LLMs in applying their vast pretrained knowledge more efficiently and aligned with human intents. This paradigm has become the central alignment pipeline. RLHF involves two main stages, with the reward model (RM) serving as a proxy for human preferences. In the first stage, the RM is trained on pairwise preference data, learning to predict the quality and alignment of candidate responses with human judgment. In the second stage, this RM scores outputs from the target policy LLM, providing optimization signals for reinforcement learning. Therefore, the quality of the reward model directly governs the effectiveness of learning for the policy LLM.
@@
However, reward modeling is inherently challenging. Human preference is often ambiguous and noisy(liu2025skyworkrewardv2scalingpreferencedata; liu2024rm; cui2023ultrafeedback). Furthermore, it tends to span a wide range of tasks with varying alignment goals. For instance, in safety-oriented scenarios, the preferred response may be refusing to answer harmful queries. In contrast, reasoning-heavy tasks such as coding depend on precise, step-by-step correctness. A reward model must therefore learn to faithfully capture these varied notions of “good” behavior across diverse data. To achieve this, it must extract equally reliable learning signals from both broad, easy-to-separate cases and fine-grained, subtle distinctions. This makes the design of training objective and its update dynamics especially important, since even minor bias in how models learn the different types of data can undermine the alignment process.
Figure 1: An illustration where a pair (top) receives inherently large update due to large representation distance, and a pair (bottom) receive weak update due to small distance
@@
The Bradley-Terry (BT) loss (bradley1952rank) has become the standard objective in pairwise reward modeling due to its simplicity and probabilistic grounding. Despite its wide adaptation (zhong2025comprehensivesurveyrewardmodels), the update dynamics under the BT loss remain under-explored. As in most supervised learning settings, we expect the magnitude of reward model update to be driven by prediction error, i.e., how wrong the model is in ranking chosen versus rejected responses. When the model misranks a pair by a large margin, updates should be strong to correct the error; when the predicted ranking is correct, updates should be weak to avoid unnecessary shifts.
@@
However, our analysis shows that this is not the case under the BT objective. By decomposing the BT gradient, we identify two factors that determine update size: (1) prediction error, calculated as the difference between the true rank and the predicted rank, and critically, (2) the representation distance between the pair of chosen and rejected responses. This coupling means that regardless of prediction accuracy, small-distance pairs tend to receive weak updates, while highly dissimilar pairs can trigger large updates even when correctly ranked. This behavior contradicts typical intuition, introducing spurious signals that undermine learning efficiency.
@@
In fact, the diversity of preference data highlights this issue in how the training loss translates pairwise comparisons into learning signals. Figure 1 illustrates this issue with two examples. In example 1, drawn from Safety task, the chosen response refuses a harmful query while the rejected response attempts to answer it. In example 2, from Reasoning task, the chosen response provides the correct coding solution while the rejected response contains a logical error that renders it entirely wrong. Both pairs reflect a clear notion of “good” versus “bad,” yet they differ significantly in representation distance: the safety pair is far apart in representation space, while the reasoning pair is close. Under the BT loss, such cases become problematic. Rather than true prediction error, representation distance disproportionately dictates the learning signal. Figure 2 shows that this effect extends beyond isolated cases and is visible across datasets.
@@
    We show that this coupling introduces spurious learning signals: small-distance pairs receive inherently weak updates even when misranked, while large-distance pairs receive strong updates. This undermines the model’s learning from fine-grained distinctions.
@@ 2 Method
In this section, we first decompose the gradient of BT-loss, and establish its connection to representation distance between response pairs. We provide evidence to illustrate how this leads to biased updates, and finally introduce NormBT as proposed solution.
@@ 2.1 Gradient Norm Analysis
For clarity, we abbreviate $r_{w}=r_{\theta}(x,y_{w}),r_{l}=r_{\theta}(x,y_{l})$, and the difference in rewards $d=r_{w}-r_{l}$. Under this notation, the gradient of equation 1 for a pair $(y_{w},y_{l})$ takes the form
@@
The gradient norm $\|\nabla_{\theta}\mathcal{L}\|$ directly reflects how much the model is updated to correct the error, which we now decompose.
@@
$\leq\big|\sigma(d)-1\big|\,{\color[rgb]{0,0.6,0}\|\mathbf{w}_{s}\|\cdot L_{g}\,\|h_{w}-h_{l}\|}.$ (5)
Secondly, at the linear score layer $\mathbf{w}_{s}$,
$\|\nabla_{\mathbf{w}_{s}}\mathcal{L}\|$ $=\big|\sigma(d)-1\big|\cdot{\color[rgb]{0.9,0.45,0}\|h_{w}-h_{l}\|}.$ (6)
Summing both contributions from $\phi,\mathbf{w}_{s}$, we have the following expression deriving from
$\|\nabla_{\theta}\mathcal{L}\|$ $=\big|\sigma(d)-1\big|\cdot\big\|\nabla_{\theta}(r_{w}-r_{l})\big\|$
$\leq\big|\sigma(d)-1\big|\cdot\sqrt{{\color[rgb]{0,0.6,0}\big(\|\mathbf{w}_{s}\|\cdot L_{g}\|h_{w}-h_{l}\|\big)^{2}}+{\color[rgb]{0.9,0.45,0}\|h_{w}-h_{l}\|^{2}}}$
$=\underbrace{\big|\sigma(d)-1\big|}_{\text{prediction error}}\cdot\underbrace{\big(k\|h_{w}-h_{l}\|\big)}_{\text{representation distance}},$ (7)
@@
This decomposition makes clear that the strength of update depends jointly on (1). prediction error, or how well the model already ranks the responses as captured by the reward difference $d$, and (2). model’s sensitivity to the pair, which in practice is governed by representation distance.
Motivations.
@@
The term $\nabla_{\theta}(r_{w}-r_{l})$ reflects how to adjust $\theta$ so that the chosen reward $r_{w}$ increases while the rejected reward $r_{l}$ decreases. It captures the sensitivity of model parameters to this pair. As we have shown through decomposing into backbone and score layer contributions in Eq. 7, this sensitivity is closely related to the representation distance $\|h_{w}-h_{l}\|$. When representations distance between $(y_{w},y_{l})$ is small, their gradients nearly cancel, producing only a small update. When the representations are far apart, the gradient difference is larger, giving a stronger update.
To better illustrate this decomposition, Figure 1 presents two representative examples from preference data. In the first example, $(y_{w},y_{l})_{\text{Safety}}$ are dissimilar responses, yielding a large representation distance term $\|\nabla_{\theta}(r_{w}-r_{l})\|$. In the second example, the two responses differ by a logical error, making them nearly identical in form and thus close in representation distance. Thus the scale factor in update magnitude of Eq. 7 is naturally larger for $(y_{w},y_{l})_{\text{Safety}}$ than that for $(y_{w},y_{l})_{\text{Reason}}$. Appendix F shows additional examples drawn from evaluations.
Figure 2: Gradient Information Across Dataset. Gradient norms (thus update sizes) vary widely by task under the BT-loss, corresponding to the variation in representation distance on the right. In particular, Reasoning pairs exhibit the smallest distance and correspondingly the weakest updates.
Q1: Does the disparity in representation distance appear across dataset?
Figure 3: Update sizes at fixed reward difference.
@@
This effect is not limited to isolated cases but extends across datasets. We illustrate this using RewardBench (lambert2024rewardbenchevaluatingrewardmodels), a benchmark for reward models that provides categorized preference data (Chat, Chat-Hard, Safety, and Reasoning), allowing easy understanding for cross-task comparisons. Figure 2 compares the distribution of gradient norms (left) and representation distance $\|\nabla_{\theta}(r_{w}-r_{l})\|$ (right) across pairs in RewardBench. The left panel shows that the gradient update sizes under the BT objective vary widely by task, with the Reasoning category exhibiting the smallest average gradient norm, despite being a domain where fine-grained distinctions are crucial. The right panel reveals why: Reasoning pairs also have the smallest representation distance on average, whereas Safety and Chat pairs are much farther apart.
@@
This observation directly supports our intuition. Since BT-loss couples prediction error with representation distance (Eq. 7), updates for small-distance pairs are inherently weaker. While such small-distance, weak-update cases appear frequently in reasoning tasks under this specific data distribution, the underlying issue is more general: it arises from the representation distance, rather than any single task domain.
Q2: What happens if representation distance outweighs prediction error?
Figure 3 provides a perspective showing how representation distance can dominate, by comparing pairs with similar prediction errors as captured by the reward difference $d$. Even within each bin, which holds prediction error nearly fixed, update size grows with representation distance.
@@
In preference learning, minor differences can decisively determine correctness. For example, a single logical mistake in reasoning tasks renders an otherwise valid solution entirely wrong. If such pairs inherently receive only weak updates due to small representation distance, the model under-corrects clear misrankings and fails to separate chosen from rejected behavior. Conversely, when certain pairs can trigger large updates regardless of accuracy, the reward model may overemphasize superficial differences. We argue that update strength should be more aligned with true prediction errors. When representation distance dominates, the learning signal may be misaligned or misleading for optimization.
@@ 2.2 Sample-Wise Normalization
To align BT update sizes more directly with prediction error, we introduce into the loss function a detached per-pair weight $w_{i}(x,y_{w},y_{l})$ that scales inversely with the norm of representation distance. In principle, this weight can be defined as
$w_{i}=\frac{1}{\|\nabla_{\theta}(r_{w}-r_{l})\|}.$ (8)
Thus the gradient norm simplifies to
$\bigl\|\nabla_{\theta}\hat{\mathcal{L}}\bigr\|=w_{i}\,\bigl|\sigma(d)-1\bigr|\,\bigl\|\nabla_{\theta}(r_{w}-r_{l})\bigr\|=\bigl|\sigma(d)-1\bigr|.$ (9)
@@
However, computing $w_{i}$ is impractical for large datasets, as it requires additional backward operations to obtain per-sample gradient information. Given the established connection $\|\nabla_{\theta}(r_{w}-r_{l})\|\propto\|h_{w}-h_{l}\|$ in Sec 2.1, we consider $\|h_{w}-h_{l}\|$ as a direct proxy, and define the pair-wise weight
$\tilde{w_{i}}=\frac{1}{\|h_{w}-h_{l}\|},$ (10)
This proxy offers three key advantages. First, it is simple and efficient: $(h_{w}-h_{l})$ can be obtained conveniently during the forward pass without additional computations. Second, it is empirically justified: beyond theoretical derivations in Eq. 7, we provide empirical evidence that this proxy is well-correlated with the full representation distance $\|\nabla_{\theta}(r_{w}-r_{l})\|$ in Appendix H. Third and importantly, this proxy is not just approximate but exact at the linear score head (Eq. 6). Since the score head directly controls the reward prediction, it is also natural to view its gradient as a convenient normalization signal.
Finally, the use of $\|h_{w}-h_{l}\|$ admits an alternative, simple interpretation. As this proxy corresponds to the final layer embedding distance between the response pair, it provides an intuitive viewpoint. Pairs with small embedding differences often correlate with responses that are e.g., close in form, structure, semantics, etc, such as the Reasoning example shown in Figure 1. From our analysis, their updates are inherently weak under the BT objective. Normalizing by $\frac{1}{\|h_{w}-h_{l})\|}$ compensates for this and upweights such pairs, allowing the training signal to allocate more attention to fine-grained distinctions. Conversely, highly dissimilar responses naturally receive large updates and are thus down-weighted, reducing the tendency to overemphasize superficial differences in representation space. Appendix C further expands on the motivation.
@@
Since the scale of embeddings can vary during training, raw values of $\tilde{w}$ can drift and tighten or loosen normalization inadvertently (Appendix G). To make the weighting $\tilde{w}$ invariant to such scale drift, we use an exponential moving average (EMA) of the batch statistics.
@@
Concretely, the NormBT objective incorporates both the proxy for representation distance and EMA-based stabilization, given by the following
$\mathcal{L}_{\text{NormBT}}(\theta)$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[{\color[rgb]{0,0,1}{\tilde{w}(y_{w},y_{l})}}\cdot\log\,\sigma\big(r_{w}-r_{l}\big)\right].$ (13)

### v2

@@ Abstract
Reward models are central to Large Language Model (LLM) alignment within the framework of RLHF. The standard objective used in reward modeling is the Bradley-Terry (BT) loss, which learns from pairwise data consisting of chosen and rejected responses. In this work, we analyze the per-sample gradient of BT-loss and show spurious learning signals due to representation distance. In particular, BT gradient norm scales with two distinct components: (1) prediction error, reflected by the difference in predicted rewards between chosen and rejected responses, and critically, (2) representation distance between the pair measured in the output space of the final layer. While the first term captures the intended training signal, the second term can significantly impact the update magnitude and misalign learning. Specifically, pairs with small representation distance often receive vanishingly weak updates, even when misranked, while pairs with large distance receive disproportionately strong updates. This leads to gradients from large-distance pairs to overshadow those from small-distance pairs, where fine-grained distinctions are especially important. To overcome this limitation, we propose NormBT, an adaptive pair-wise normalization scheme that rescales update to balance representation-driven effects and focuses learning signals on prediction error. NormBT is a lightweight, drop-in modification to BT loss with negligible overhead. Across various LLM backbones and datasets, NormBT improves reward model performance consistently, with notable gains of over 5% on the Reasoning category of RewardBench, which contains numerous fine-grained pairs.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as the leading framework for aligning Large Language Models (LLMs) with human preferences (Christiano et al., 2017; Ziegler et al., 2020; Ouyang et al., 2022). Instead of traditional supervision, RLHF uses preference signals to guide LLMs in applying their vast pretrained knowledge more efficiently and aligned with human intents. This paradigm has become the central alignment pipeline. RLHF involves two main stages, with the reward model (RM) serving as a proxy for human preferences. In the first stage, the RM is trained on pairwise preference data, learning to predict the quality and alignment of candidate responses with human judgment. In the second stage, this RM scores outputs from the target policy LLM, providing optimization signals for reinforcement learning. Therefore, the quality of the reward model directly governs the effectiveness of learning for the policy LLM.
@@
However, reward modeling is inherently challenging. Human preference is often ambiguous and noisy (Liu et al., 2025, 2024c; Cui et al., 2023). Furthermore, it tends to span a wide range of tasks with varying alignment goals. For instance, in safety-oriented scenarios, the preferred response may be refusing to answer harmful queries. In contrast, reasoning-heavy tasks such as coding depend on precise, step-by-step correctness. A reward model must therefore learn to faithfully capture these varied notions of “good” behavior across diverse data. To achieve this, it must extract equally reliable learning signals from both broad, easy-to-separate cases and fine-grained, subtle distinctions. This makes the design of training objectives and their update dynamics especially important, since even minor bias in how models learn the different types of data can undermine the alignment process.
Figure 1: Representation-distance bias in BT gradients. Under BT objective, update magnitudes scale with representation distance: “far-apart” pairs (top) can produce inherently large gradients, while “close” pairs (bottom) produce weak gradients, even when the model is equally wrong in its predictions for both pairs. Our method NormBT introduces a per-pair scale factor, normalizing the gradient contributions by representation distance, such that update magnitudes depend primarily on prediction errors.
@@
The Bradley-Terry (BT) loss (Bradley and Terry, 1952) has become the standard objective in pairwise reward modeling due to its simplicity and probabilistic grounding. Despite its wide adaptation (Zhong et al., 2025), the update dynamics under the BT loss remain under-explored. As in most supervised learning settings, we expect the magnitude of reward model update to be driven by prediction error, i.e., how wrong the model is in ranking chosen versus rejected responses: when the model misranks a pair by a large margin, updates should be strong to correct the error; when the predicted ranking is correct, updates should be weak to avoid unnecessary shifts.
@@
However, our analysis shows that this is not the case under the BT objective. We decouple the two factors that determine BT update size: (1) prediction error, calculated as the predicted reward difference between chosen and rejected response, and critically, (2) the representation distance between the pair. This coupling means that regardless of prediction accuracy, small-distance pairs tend to receive weak updates, while large-distance pairs naturally receive large updates even when correctly ranked (Figure 2(a)). This behavior contradicts typical intuition, introducing signals that undermine learning in reward models.
@@
This issue becomes especially consequential in realistic preference datasets, which contains meaningful comparisons where some pairs could differ dramatically in content, while others are nearly identical except for a single decisive mistake in response. Figure 1 shows two example pairs. In example 1, drawn from Safety task, the chosen response refuses a harmful query while the rejected response answers it. In example 2, from Reasoning task, the chosen response provides the correct coding solution while the rejected response contains a logical error that renders it entirely wrong. Both pairs reflect a clear notion of “good” versus “bad,” yet they differ significantly in representation distance: the safety pair is far apart in representation space, whereas the model struggles to separate the reasoning pair and therefore assigns them highly similar representations. Under the BT loss, such cases become problematic. Rather than true prediction error, representation distance dictates the learning signal. Figure 2(b) shows that this effect extends beyond isolated cases and is visible across datasets.
@@
    We show that this coupling misaligns learning signals: small-distance pairs receive inherently weak updates even when misranked, while large-distance pairs receive strong updates. This undermines the model’s learning from fine-grained distinctions.
@@ 2 Method
In this section, we first decompose the gradient of BT-loss, and establish its connection to the representation distance between response preference pairs. We provide evidence to illustrate how this leads to biased updates, and finally introduce NormBT as the proposed solution.
@@ 2.1 Gradient Norm Analysis
For clarity, we abbreviate $r_{w}=r_{\theta}(x,y_{w}),r_{l}=r_{\theta}(x,y_{l})$, and difference in rewards $d=r_{w}-r_{l}$. Under this notation, the gradient of Eq.1 for a pair $(y_{w},y_{l})$ takes the form
@@
The gradient norm $\|\nabla_{\theta}\mathcal{L}\|$ directly reflects how much the model is updated, which we now decompose.
@@
$\leq\big|\sigma(d)-1\big|\,{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\|\mathbf{w}_{s}\|\cdot L_{g}\,\|h_{w}-h_{l}\|}.$ (5)
Secondly, at the component of score layer $\mathbf{w}_{s}$,
$\|\nabla_{\mathbf{w}_{s}}\mathcal{L}\|$ $=\big|\sigma(d)-1\big|\cdot{\color[rgb]{0.9,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.45,0}\|h_{w}-h_{l}\|}.$ (6)
Summing both contributions from $\phi,\mathbf{w}_{s}$, we have the following expression for the norm of Eq.1:
$\|\nabla_{\theta}\mathcal{L}\|$ $=\big|\sigma(d)-1\big|\,\big\|\nabla_{\theta}(r_{w}-r_{l})\big\|$
$\leq\big|\sigma(d)-1\big|\,\sqrt{\begin{aligned} &{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\big(\|\mathbf{w}_{s}\|\,L_{g}\,\|h_{w}-h_{l}\|\big)^{2}}\\ &\quad+{\color[rgb]{0.9,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.45,0}\|h_{w}-h_{l}\|^{2}}\end{aligned}}$
$=\underbrace{\big|\sigma(d)-1\big|}_{\mathclap{\text{prediction error}}}\cdot\underbrace{\big(k\|h_{w}-h_{l}\|\big)}_{\mathclap{\text{representation distance}}},$ (7)
@@
This decomposition makes clear that the strength of update depends jointly on (1). prediction error, i.e., how well the model currently ranks the responses, as captured by the predicted reward difference $d$, and (2). model’s sensitivity to the pair, which by derivation in Eq. 7 is governed by representation distance.
Interpretations.
@@
The term $\nabla_{\theta}(r_{w}-r_{l})$ reflects how $\theta$ is adjusted so that the chosen reward $r_{w}$ increases while the rejected reward $r_{l}$ decreases. It captures the sensitivity of model parameters to this pair. As shown by decomposing into backbone $\phi$ and score layer $\mathbf{w}_{s}$ contributions in Eq. 7, this sensitivity is closely related to the representation distance $\|h_{w}-h_{l}\|$. Consequently, for two pairs with the same prediction error (i.e., the same $\big|\sigma(d)-1\big|$ term in the BT gradient), the pair with small representation distance $\|h_{w}-h_{l}\|$ gives nearly canceling gradients and receives only a weak update, whereas the pair with a large distance receives a stronger update.
To show this concretely, Figure 1 presents two preference pairs with very different representation distances. In the first pair, $(y_{w},y_{l})_{\text{Safety}}$ are semantically and stylistically dissimilar, leading to a large representation distance term $\|h_{w}-h_{l}\|$. In the second pair, $(y_{w},y_{l})_{\text{Reason}}$ differ by a logical error and are nearly identical in surface form. While humans would prefer the correct $y_{w}$ over the wrong $y_{l}$, this pair is harder for the model to separate, producing a smaller representation distance. As a result, Eq. 7 implies that the BT update magnitude for $(y_{w},y_{l})_{\text{Reason}}$ is naturally weaker than that for $(y_{w},y_{l})_{\text{Safety}}$, systematically under-training pairs where the model struggles to distinguish. Appendix D presents additional examples drawn from evaluations.
(a) Gradient norms, binned by reward diff. $d$
(b) Gradient Norm by Category
(c) Representation Distance by Category
Figure 2: Gradient Analysis Across Dataset. (a) Even for pairs with the same prediction error $d$ (i.e., pairs within each color group), its BT gradient norm, hence update magnitude, still scales strongly with representation distance $\|h_{w}-h_{l}\|$. (b) BT gradient norms vary substantially across RewardBench categories, with Reasoning receiving the weakest updates. (c) This disparity in (b) aligns with category-wise representation distances: Reasoning pairs are also closest in representation space, while others are farther apart.
Quantitative analysis 1: How does representation distance “dominate” update magnitudes?
Figure 2(a) further isolates the effect of representation distance by controlling for prediction error: we group preference pairs into bins with nearly identical error term $\big|\sigma(d)-1\big|$ (equivalently, similar $d=r_{w}-r_{l}$). Within each bin, which holds prediction error term nearly fixed, the observed gradient norms increase systematically with representation distance. This provides direct evidence showing that the distance term dominates the learning signal even when the model is “equally wrong.”
Quantitative analysis 2: How prevalent is this representation distance disparity in preference data?
@@
This effect is not limited to isolated cases but extends across datasets. We illustrate this using RewardBench (Lambert et al., 2024), a benchmark for reward models that provides categorized preference data (Chat, Chat-Hard, Safety, and Reasoning), allowing easy understanding for cross-task comparisons. Figure 2(b) compares the distribution of gradient norms and  2(c) compares $\|\nabla_{\theta}(r_{w}-r_{l})\|$ across pairs in RewardBench. The left panel shows that the gradient norms (hence update sizes) under the BT objective vary widely by task. In particular, the Reasoning category exhibits the smallest average gradient norm, despite being a domain where fine-grained distinctions are crucial. The right panel reveals why: Reasoning pairs also have the smallest representation distance, whereas e.g., Safety and Chat pairs are much farther apart in representation space.
@@
This observation provides dataset-wise evidence for our interpretations. Since BT-loss couples prediction error with representation distance (Eq. 7), updates for small-distance pairs are inherently weaker. Nonetheless, we note that while such small-distance, weak-update cases appear frequently in reasoning tasks under this specific data distribution, the underlying issue is more general: it arises from the representation distance, rather than any single task domain.
Motivation.
@@
In preference learning, minor differences can decisively determine correctness. For example, a single logical mistake in reasoning tasks can make an otherwise valid solution entirely wrong. In well-curated preference datasets, small representation distance often reflects meaningful and hard pairs that the current model fails to separate despite a clear human preference (e.g., correct vs. subtly flawed code). If such pairs inherently receive only weak updates due to small representation distance $\|h_{w}-h_{l}\|$, the model under-corrects clear misrankings and receives the least training signal precisely where it struggles most. Conversely, when certain pairs can produce large updates regardless of accuracy, the reward model may overemphasize superficial differences. These observations motivate aligning reward model update strength with prediction error rather than representation distance.
@@ 2.2 Sample-Wise Normalization
To align BT update sizes directly with prediction error, we introduce into the loss function a per-pair weight $w_{i}(x,y_{w},y_{l})$ that scales inversely with representation distance. In principle, this weight can be defined as
$w_{i}=\text{sg}\Big(\frac{1}{\|\nabla_{\theta}(r_{w}-r_{l})\|}\Big),$ (8)
where $\text{sg}(\cdot)$ denotes the stop-gradient operator. Thus the gradient norm simplifies to
$\bigl\|\nabla_{\theta}\hat{\mathcal{L}}\bigr\|={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}w_{i}}\,\bigl|\sigma(d)-1\bigr|\,\bigl\|\nabla_{\theta}(r_{w}-r_{l})\bigr\|=\bigl|\sigma(d)-1\bigr|.$ (9)
@@
Computing $w_{i}$ is impractical at scale, as it requires additional backward operations to obtain per-sample gradient information. Instead, analysis in Sec 2.1 establishes that it is well-approximated by the final-layer representation distance: $\|\nabla_{\theta}(r_{w}-r_{l})\|\propto\|h_{w}-h_{l}\|$ (Eq. 7). We therefore consider $\|h_{w}-h_{l}\|$ as a direct proxy, and define the pair-wise weight
$\tilde{w_{i}}=\text{sg}\Big(\frac{1}{\|h_{w}-h_{l}\|}\Big),$ (10)
This proxy offers three key advantages: (1) Efficient. $(h_{w}-h_{l})$ is available from the forward pass without additional computations. (2) Well-founded. This proxy is not just approximate but exact at the linear score head (Eq. 6). Since the score head directly produces the reward prediction, it is natural to view its gradient as a normalization signal. (3) Admits an intuitive interpretation. $\|h_{w}-h_{l}\|$ corresponds to the final-layer representation distance between the response pair. Small representation gaps often arise when responses are close in surface form or semantics (e.g., the Reasoning pair in Figure 1), which are precisely the fine-grained distinctions that are hardest for the model to separate. Normalizing by $\frac{1}{\|h_{w}-h_{l})\|}$ prevents further suppressing the update strength by upweighting such subtle but decisive pairs. Conversely, highly dissimilar responses that naturally induce large BT updates and are therefore downweighted, reducing the tendency to overemphasize superficial differences in representation space.
The correlation between this proxy and full representation distance $\|\nabla_{\theta}(r_{w}-r_{l})\|$ is shown in Appendix F.
@@
Since the scale of embeddings can vary during training, raw values of $\tilde{w}$ may drift and tighten or loosen normalization inadvertently (Appendix G). To make $\tilde{w}$ invariant to such scale drift, we use an exponential moving average (EMA) of the batch statistics.
@@
Concretely, the NormBT objective incorporates both the proxy for representation distance and EMA-based stabilization. For a preference sample $s=(x,y_{c},y_{r})$, it is given by
$\mathcal{L}_{\text{NormBT}}(\theta)$ $=-\mathbb{E}_{s\sim D}\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\tilde{w}(y_{w},y_{l})}}\cdot\log\,\sigma\big(r_{w}-r_{l}\big)\right].$ (13)